# DeToNATION: Decoupled Torch Network-Aware Training on Interlinked Online Nodes

### v3

@@ Conclusion
Appendix B B: DeMo Hyperparameters
@@
The sign-function [^3] https://docs.pytorch.org/docs/2.7/generated/torch.sign.html introduces great reduction in required data transferred, reducing the actual communication to a ternary system with minimal information exchange. This design-choice was introduced in DeMo (Peng, Quesnelle, and Kingma 2024), however, not justified. We report loss on the validation set in Figure 9. Interestingly, employing the sign-function to gradients before synchronization, generally yields significant performance-gains with a huge gap in the validation-results with DeMo and DiLoCo leading performance, closely followed by the Random scheme, with Striding under-performing quite significantly.
@@
Average Time per Optimization Step
We analyze the average time per step for $10$, $100$, $1000$ and 10.00010{.}000 Mbps network connections training the T5-Large and ViT-B. The experiment is conducted on a controlled network between two local dedicated nodes. We compare DeMo and Random replication schemes with underlying DeMo-SGD optimizer with compression-rates $\frac{1}{16}$ and $\frac{1}{32}$, against the Decoupled-AdamW optimizer (ours). We report our results in Figure 10. As expected, as the bandwidth decreases, the importance of the compression-rate becomes more pronounced, where the compression of $\frac{1}{32}$ clearly demonstrates to be fastest followed by $\frac{1}{16}$ with $\frac{1}{1}$, being last, dictated by the amount of data to communicate, with Random DeMo-SGD being the fastest, as demonstrated both in Figure 10(a) and 10(b). This is especially significant at bandwidths lower than 500Mbps, which is a level of bandwidths comparable to those available in HPCs, in practice. Random replication is approx. $3.33$ times faster than DeMo-replication at 1010Mbps, while being approx. $18$ times faster than Decoupled-AdamW with full replication (Figure 10(b)). Generally we see that Random replication with compression $\frac{1}{16}$ scales similarly to DeMo replication with compression $\frac{1}{32}$, demonstrating exactly that Random replication is around twice as fast as DeMo, as expected from DeMo transferring twice the amount of data, at the same compression rate.
(a) T5-Large on 2 nodes across employing different replicators using DeMo-SGD and Decoupled AdamW as base optimizers
(b) ViT-B on 2 nodes across employing different replicators using DeMo-SGD and Decoupled AdamW as base optimizers
Figure 10: Average time per optimizer step for T5-Large and ViT-B in bandwidth-constrained setups.
@@
The most significant momenta are shared in full precision by previous work (Peng, Quesnelle, and Kingma 2024). Naturally, we need to confirm the necessity to transfer double precision on the cost of doubled data footprint, visualized in Figure 13. We validate this choice in Figure 14, where we confirm DeMo and full precision is superior, followed close by full sync (DiLoCo) full-precision. Interestingly, the full-sync scheme is mostly unaffected by the data precision, whereas both DeMo and Random replication change significantly with precision, where higher precision, leads to better performance.
@@
Appendix C C: T5-Large Train Losses
In Figure 15 we report the train-losses for the T5 translation task experiments in the main paper.
@@ Appendix C C: T5-Large Train Losses
Appendix D D: ViT Train Losses
In Figure 16 we report the train-losses for the ViT image classification experiments in the main paper.
@@ Appendix D D: ViT Train Losses
Appendix E E: Ethical Considerations
@@
Appendix F Reproducibility
[…]

### v4

@@ Appendix B Reproducibility Checklist
Appendix D B: DeMo Hyperparameters
Here, we study hyperparameters for DeMo replication, including sign, TopK, transfer-dtype. The experiments are conducted with Hybrid-FSDP on two nodes with two accelerators each. We run training for 20 epochs without weight decay. The optimizer and the replication scheme depends on the specific experiment.
@@ Appendix D B: DeMo Hyperparameters
The sign-function [^3] https://docs.pytorch.org/docs/2.7/generated/torch.sign.html introduces great reduction in required data transferred, reducing the actual communication to a ternary system with minimal information exchange. This design-choice was introduced in DeMo (peng2024demodecoupledmomentumoptimization), however, not justified. We report loss on the validation set in Figure 10. Interestingly, employing the sign-function to gradients before synchronization generally yields substantial performance gains with a large margin in the validation results, with DeMo and DiLoCo yielding the strongest performance, closely followed by Random, while Striding is underperforming.
@@
The most significant momenta are shared in full precision by previous work (peng2024demodecoupledmomentumoptimization). Naturally, we need to confirm the necessity to transfer double precision on the cost of doubled data footprint, visualized in Figure 13. We validate this choice in Figure 14, where we confirm DeMo and full precision is superior, followed close by full sync (DiLoCo) full-precision. Interestingly, the full-sync scheme is mostly unaffected by the data precision, whereas both DeMo and Random replication change significantly with precision, where higher precision, leads to better performance.
@@
Appendix E C: Experiment Train Losses
Here, we report train-loss curves for the experiments in the main paper. In Figure 15 we report the train-losses for the T5 translation task experiments, and in Figure 16 we report the train-losses for the ViT image classification experiments. Figure 17 shows the non-zoomed training loss of the OLMo2 experiment in the main paper (Figure 3.
@@ Appendix E C: Experiment Train Losses
Figure 17: OLMo2 1B train loss over 10K training steps on Dolma v1.6 using different replicators and compression rates. All experiments, except the Hybrid-FSDP baseline with AdamW on 2 nodes, use DeMo-SGD.
Appendix F D: Average Time per Optimizer Step
Figure 18 shows the average time per optimizer step for T5-Large on 2 nodes across different replicators and bandwidth limits.
Figure 18: Average time per optimizer step for T5-Large on 2 nodes across employing different replicators using DeMo-SGD and Decoupled AdamW as base optimizers
Appendix G E: Ethical Considerations
@@
Appendix H F: SGD vs. Decoupled AdamW
AdamW (kingma2014adam) is de-facto default optimizer, having demonstrated superior performance across many domains and tasks. However DeMo (peng2024demodecoupledmomentumoptimization) is employing SGD (which we denote DeMo-SGD, we differentiate as it accumulates momenta.) as underlying optimizer. To investigate this design-choice and broaden the investigation, we set out to compare DeMo-SGD and our Decoupled-AdamW with employing DeMo, DiLoCo (douillard2023diloco), Random, and Striding replication schemes. We report validation loss in Figure 19. The experiments are conducted such that the bandwidth usage is identical in each setting (we denote this as the compression-rate), as it would be defined by hardware in practice. DeMo-SGD with Random replication demonstrates best performance. DeMo-SGD yields clearly better at generalization to the validation-set, with Random replication being best. Decoupled-AdamW and DeMo-SGD with DiLoCo replication following as seconds and third best, respectively. DeMo-SGD is improving overall as fast as the DiLoCo variants, but overall worse. Decoupled-AdamW with DeMo and Random replication schemes yields worst performance.